### v2

@@ 3.4 Training and Inference
During training, we focus solely on optimizing the style controller, leaving the parameters of the pre-trained diffusion model unaltered. The style controller is trained using a dataset of paired images and text. Still, it can train without text prompts, as only image prompts effectively guides the final generation. The training objective remains consistent with that of the original SD:
@@
During the training phase, we consistently employ random omission of image conditions to facilitate classifier-free guidance during inference.
$\begin{split}\hat{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)=w\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)+(1-w)\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t),\end{split}$ (4)
If the image condition is omitted, we replace the CLIP image embedding with a zero vector.
@@ 4.1 Datasets
Figure 3: In this study, data processing involved the following four steps: (1) Collecting raw data, including images and text sources. (2) Cleaning the data to eliminate noise, duplicates, and irrelevant information. (3) Preprocessing of data, such as resizing images and converting formats. (4) Creating matching pairs of text and images for model training.
Limited datasets are available for landscape paintings. The traditional landscape painting dataset collected by Xue  Xue (2021) holds significant value. Building upon this dataset, we have expanded our collection to include various styles of landscape paintings and standardized the corresponding text descriptions. This effort aims to contribute positively to the research on landscape painting generation, as illustrated in Figure 3.
Collection: Our investigation revealed a scarcity of landscape painting datasets available for research. Currently, the Traditional Landscape Painting Dataset Xue (2021) is the only accessible dataset, but it has limitations in terms of data volume and lacks categorization or identification of the image data. To supplement this, we have curated a collection of 500 digital paintings sourced through websites such as Google and Baidu, artwork websites, photographs from landscape painting books, and digital museum databases. The size distribution of these samples is detailed in Table 1 below. To ensure representativeness, we included works from various periods and historical stages.
Table 1: This table is based on the minimum images table for statistical information. Images with small edges greater than or equal to 512 pixels are our effective data and can be used for model training in the future.
Category Count
Less than 512 319
Equal to 512 2617
Greater than or equal to 512 16
@@
Cleaning: We enlisted the help of professors and landscape painters to assist in the image-cleaning process. To ensure the dataset’s quality, these experts manually removed paintings that did not depict landscapes and those with dimensions smaller than 512 pixels or with unclear imagery. Landscape painting is a comprehensive art form that integrates poetry, calligraphy, seal cutting, and painting, unique aspects of this artistic tradition. Directly using these paintings as training data can result in the model generating images with inexplicable elements or words. To better emphasize the theme and beauty of landscape paintings while removing the interference of calligraphy and seal engraving, we utilized image processing software like Adobe Photoshop to repair the damaged area, enhancing image clarity and quality. We adjusted the brightness, contrast, hue, and saturation parameters to make the images sharper and brighter and to highlight their details and features. In addition, we manually used the repair tool to meticulously restore the integrity and beauty of the original images, employing multiple overlapping repair pen strokes to achieve a more complete restoration.
@@
Preprocessing: Since the training data for this study should be in $512\times 512$ format, I processed the images accordingly. First, each image was scaled so that its shorter side was 512 pixels, maintaining the original aspect ratio. For images with an aspect ratio less than 1.5, a $512\times 512$ section was cropped from the center and saved. For images with an aspect ratio greater than 1.5, after cropping the $512\times 512$ section from the center, the center point was moved 256 pixels in both directions along the longer side to crop and save two additional images. If the cropped image was not square, the process was halted, and the current cropped image was discarded.
@@
Text-image Pair Generation: We enlisted landscape painters and professors to categorize the previously processed photos and the Traditional Landscape Painting Dataset. The categories identified are azure green landscape, golden splendor landscape, ink wash landscape, and light vermilion landscape. To facilitate successful model training, the data format for this experiment is image-text pairs. Although image-text pairs previously required manual annotation, we now utilize BLIP-2 to automatically generate the corresponding text information for the collected images. Although the text generated by BLIP-2  Li et al. (2022) may be inaccurate, we invited professional teachers to refine and augment the text descriptions specific to landscape paintings. To increase the diversity of text descriptions, we provided multiple expressions for each text and specified the painting type within the descriptions. Finally, we saved the image paths and corresponding text descriptions in JSON files. This newly curated dataset, comprising the base data and the additional image and image-text pairs, was used to train the new model.
@@ 4.2 Evaluation Metrics
Learned Perceptual Image Patch Similarity: LPIPS is a metric to evaluate image similarity. It utilizes feature representations of deep learning models to measure the perceptual similarity between two images. Unlike traditional pixel-based similarity metrics (e.g., mean square error or peak signal-to-noise ratio), LPIPS focuses more on the perceptual quality of the image, i.e., the perception of image similarity by the human visual system. The formula for LPIPS is given by:
@@
Gram Matrix: The gram matrix is a metric for evaluating the stylistic similarity of images and is commonly used in style migration tasks. It captures the texture information and stylistic features of an image by computing the inner product between the feature maps of a convolutional neural network. Specifically, the gram matrix describes the correlation between different channels in the feature map, thus reflecting the overall texture structure of the image.
@@
Histogram Similarity (Bhattacharyya Distance): The histogram similarity is a metric for evaluating the similarity of color distributions of two images and is widely used in image processing and computer vision. By comparing the color histograms of the images, the similarity of the images in terms of color can be determined.The bhattacharyya Distance  Bhattacharyya (1943) is a commonly used histogram similarity metric, especially for similarity calculation of probability distributions.
@@
Chamfer Match Score: The chamfer match score is a metric for evaluating the similarity of two sets of point clouds, widely used in computer vision and graphics. It quantifies the shape similarity between two sets of point clouds by calculating the average nearest distance between them. Specifically, chamfer match score is a variant of Chamfer Distance and is commonly used for tasks such as 3D shape matching, image alignment and contour alignment.
@@
{adjustwidth}
[…]
Figure 4: This figure showcases an example of landscape painting generated using our method. Our approach is highly responsive to the canny edge maps and style reference images. Columns (a), (b), (c), and (d) represent landscape paintings generated in the same style but with different canny edge maps. Conversely, rows (e), (f), (g), and (h) depict landscape paintings generated with the same canny edge map but using different style references.
Hausdorff Distance: The hausdorff distance is a metric for evaluating the maximum distance between two sets of point sets (point clouds, contours, etc.) and is widely used in computer vision, graphics, and pattern recognition. It measures the distance between the farthest corresponding points in two point sets, reflecting the degree to which they differ geometrically.
@@
Contour Match Score: The contour match score evaluates the similarity between the contour shapes of two images, making it particularly useful in applications where the overall shape and structure of objects are crucial. This score is determined by comparing the shape descriptors of the contours in the images.
@@
4.3 Experimental Results
As shown in Figure 4, we use the LPen model to generate landscape paintings. This model learns the characteristics and styles of numerous classic landscape paintings and automatically creates new, highly realistic works. The generated landscape paintings feature typical natural elements such as mountains, rivers, trees, and stones and retain the ink color and brush techniques characteristic of traditional landscape paintings. Additionally, the LPen model can generate new works through innovation and variation.
The LPen model effectively and accurately controls image generation through edge detection and stylized reference images. Using the canny edge detection method, the LPen model extracts edge information from the input image, preserving and improving structural features during the generation process. Meanwhile, the style reference image provides detailed information about the target style, enabling the LPen model to learn and apply these features, thereby aligning the generated image with the expected visual effect. By combining these two control methods, the LPen model generates highly realistic landscape paintings and flexibly adjusts the generated images’ style and details, achieving precise control and innovative performance in traditional art creation.
@@ 4.4 Qualitative Analysis
To rigorously assess the effectiveness of the proposed LPGen model in generating high-quality landscape paintings, we conduct a comprehensive comparative analysis against several benchmark methods, including Reference Only Zhang et al. (2023), Double ControlNet Zhang et al. (2023), and Lora Hu et al. (2021).
The comparison experiment with other methods for generating landscape paintings is shown in Figure 5. In this figure, Figure 6a displays the constraint canny edge map, while Figure 6b shows the target ink-style reference image. Figure 6c presents the ink-style image generated by the Lora method. Here, the generated image exhibits noticeable noise and incorrect image information and is limited to Lora’s inherent style, failing to replicate the reference image’s style. Figure 6d illustrates the image produced by the Reference Only method. This image has a more precise contour structure but introduces redundant contours, lacks distinct style characteristics, and suffers from color inaccuracies. Figure 6e shows the ink-style image generated by the Double ControlNet method. Although the constraint contours are well-defined, the style characteristics of the reference image are not effectively learned.Finally, Figure 6f displays the images generated by LPGen.
In summary, the proposed LPGen effectively addresses significant issues such as detail loss, poor style transfer, and blurred lines in generated images, resulting in high-quality landscape paintings. As illustrated in Figure 5, our model consistently outperforms other state-of-the-art methods. LPGen not only preserves the layout and content of the target photos but also successfully captures the essence of the ink wash style, thereby achieving superior overall quality and fidelity.
@@
Figure 5: Samples of landscape paintings generated using different methods are shown below. Column (a) displays the canny edge image. Column (b) shows the reference image. Column (c) presents the image generated by Lora. Column (d) features the image generated by Reference Only. Column (e) illustrates the image generated by Double ControlNet. Finally, column (f) displays the image generated by our LPGen model.
4.5 Quantitative analysis
The results depicted in Figure 5 align with the quantitative metrics presented in Table 2. The LPIPS, gram matrix, and histogram similarity metrics provide robust and complementary measures to evaluate image similarity. LPIPS assesses perceptual similarity by leveraging deep feature representations, while the gram matrix evaluates style similarity by capturing texture correlations. Histogram similarity, on the other hand, compares color distributions by analyzing histograms. Together, these metrics form a comprehensive framework for determining how closely generated images resemble reference images in both structure and style. Additionally, the chamfer match score, hausdorff distance, and contour match score offer robust and complementary measures for evaluating edge and contour similarity. The chamfer match score focuses on the average similarity of edges, the hausdorff distance emphasizes the worst-case edge similarity, and the contour match score evaluates overall shape similarity. Collectively, these metrics provide a detailed framework for assessing how closely generated images match reference images in terms of edges and contours.
Table 2: The table compares style and structure metrics for images generated by different models. LPIPS evaluates perceptual similarity, GM (gram matrix) measures texture correlations for style similarity, and HS (histogram similarity) assesses color distribution. CMS i(chamfer match score) focuses on average edge similarity, HD (hausdorff distance) emphasizes worst-case edge similarity, and CMS ii(contour match score) evaluates overall shape similarity.
{adjustwidth}
[…]
@@ 4.5 Quantitative analysis
The first three metrics, LPIPS, the gram matrix, and histogram similarity, are utilized to analyze the structural similarity between the model outputs and the reference images. LPIPS (Learned Perceptual Image Patch Similarity) measures perceptual similarity between images, where lower LPIPS values indicate higher similarity in terms of human visual perception. According to the table, Lora (0.52) performs best in terms of LPIPS, indicating that Lora’s generated images are most perceptually similar to the reference image. The gram matrix metric measures style similarity, particularly in texture, with lower values indicating higher style similarity. The table shows that LPGen (3.40e-06) performs best regarding the gram matrix, suggesting that LPGen’s generated images have the most similar texture style to the reference image. Histogram similarity, measured using the bhattacharyya distance, evaluates the similarity of color distributions, with lower values indicating more identical color distributions. LPGen (0.72) performs best regarding histogram similarity, suggesting that LPGen-generated images have the most similar color distribution to the reference image. For structural similarity, LPGen excels in both texture style (gram matrix) and color distribution (histogram similarity), while Lora performs best in perceptual similarity (LPIPS) due to the structure controller. However, Lora requires repetitive fine-tuning and is computationally intensive. Taking these factors into account, LPGen is the best model for overall structural similarity.
@@
The last three metrics, the chamfer match score, hausdorff distance, and contour match score, are used to analyze the style similarity between different model outputs and the reference image, mainly focusing on edges and contours. The chamfer match score measures edge similarity, where higher scores indicate more remarkable similarity. LPGen (-0.15) performs best with respect to the chamfer match score, suggesting that LPGen’s generated images have the most similar edges to the reference image. The hausdorff distance measures the maximum edge distance between two images, with lower values indicating more remarkable similarity. LPGen (154.54) performs best regarding hausdorff distance, suggesting that LPGen’s generated images are closest to the reference image regarding edge similarity. The contour match score measures the similarity of contour shapes, with lower values indicating more remarkable similarity. LPGen (3.24) performs best regarding the contour match score, indicating that LPGen’s generated images have the most similar contour shapes to the reference image. LPGen outperforms other models in style similarity in all three metrics: chamfer match score, hausdorff distance, and contour match score. This demonstrates that LPGen’s generated images have the most similar edges and contour shapes to the reference image, making it the best model for style similarity.
4.6 Visual Turing Tests
In the user study, we evaluated four critical aspects of the generated images: aesthetic appeal, style consistency, creativity, and detail quality. Aesthetic Appeal: This metric evaluates the overall visual attractiveness of the images. The participants rated the images based on how pleasing they found them, considering factors such as color harmony, composition, and the emotional response evoked by the artwork. Style Consistency: This aspect examines how well the generated images adhere to a specific artistic style. Participants evaluated whether the images consistently incorporated stylistic elements of traditional landscape paintings, such as brushstroke techniques, use of space, and traditional motifs. Creativity: This aspect measures the originality and innovation of the generated images. The participants rated the images based on the novelty and inventiveness of the compositions and interpretations within the confines of traditional landscape painting. Detail Quality: This metric focuses on the precision and clarity of the finer details within the images. Participants evaluated the quality of intricate elements such as textures, line work, and depiction of natural features such as trees, rocks, and water.

### v3

@@ 3.4 Training and Inference
During training, we focus solely on optimizing the style controller, leaving the parameters of the pre-trained diffusion model unchanged. The style controller is trained using a dataset of paired images and text. Still, it can train without text prompts, as only image prompts effectively guides the final generation. The training objective remains consistent with that of the original SD:
@@
During the training stage, we consistently employ the random omission of image conditions to facilitate classifier-free guidance during inference. If the image condition is omitted, we replace the CLIP image embedding with a zero vector.
$\begin{split}\hat{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)=w\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)+(1-w)\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t).\end{split}$ (4)
@@ 3.5 Datasets
Figure 3: Experimental dataset processing workflow. The workflow begins with collecting raw image data, followed by cleaning the data to eliminate noise, duplicates, and irrelevant information. The data is then pre-processed by resizing images and converting formats. Finally, matching pairs of text and images are created for model training.
Text-to-image generative diffusion models have been slow to develop in landscape painting generation due to the lack of image-text paired datasets describing style and content. Thus, the diffusion model cannot stably generate landscape painting for specified decoration style and structure. To this end, it is imminent to establish a new dataset of interior design decorating styles. To this end, this study first constructed a dataset, LPD-3, with descriptions of styles and content, which were collected by from websites, as illustrated in Figure 3. We have expanded our collection to include various styles of landscape painting and standardized the corresponding text descriptions. This effort aims to contribute positively to the research on landscape painting generation.
Data Collection. Datasets serve as the foundation for the rapid advancement of artificial intelligence, with a few open datasets significantly contributing to this progress. For example, the traditional landscape painting dataset  Xue (2021) is the only accessible dataset, but it has limitations in terms of data volume and lacks categorization or identification of the image data. To address this problem, we firstly have curated a collection of digital paintings sourced through websites such as Baidu, artwork websites, photographs from landscape painting books, and digital museum databases. In the following, we enlisted the help of professors and landscape artists to assist in the classification process. All the landscape paintings are classified into three main categories: azure green landscape, ink wash landscape, and light vermilion landscape.
@@
Data Cleaning. To ensure the quality of the dataset, these experts manually removed paintings that did not depict landscapes and those with dimensions smaller than 512 pixels or with unclear imagery. If the collected images are used directly for training, the resulting landscape paintings may contain inexplicable elements or words. To better emphasize the theme and beauty of landscape paintings while removing the interference of calligraphy and seal cutting, we utilized image processing software like Adobe Photoshop to repair the damaged area, enhancing image clarity and quality. We adjusted the brightness, contrast, hue, and saturation parameters to make the images sharper and brighter and to highlight their details and features.
@@
Data Preprocessing. Since the training data for this study should be in $512\times 512$ format, preprocessing the cleaned data is an essential step. First, each image was scaled so that its shorter side was 512 pixels, maintaining the original aspect ratio. For images with an aspect ratio less than 1.5, a $512\times 512$ section was cropped from the center and saved. For images with an aspect ratio greater than 1.5, after cropping the $512\times 512$ section from the center, the center point was moved 256 pixels in both directions along the longer side to crop and save two additional images. If the cropped image was not square, the process was halted, and the current cropped image was discarded. The distribution of the number of pictures is shown in Table 1.
Table 1: Landscape painting type distribution of three distinct styles in our collected dataset. The data comes from Harvard, Smithsonian, Metropolitan, Baidu, and Princeton.
Harvard Smithsonian Metropolitan Baidu Princeton
Azure Green Landscape 12 192 19 181 18
Light Vermilion Landscape 18 742 256 104 273
Ink Wash Landscape 68 342 119 20 52
Total 98 1276 394 305 343
@@
Data Pairing. The data format for this experiment is image-text pairs because it allows the model to learn the association between visual content and descriptive language, enhancing its ability to generate images that align with specific textual descriptions, thereby improving the overall accuracy and relevance of the generated images. Although image-text pairs previously required manual annotation, we now utilize BLIP-2 to automatically generate the corresponding text information for the collected images. Although the text generated by BLIP-2  Li et al. (2022) may be inaccurate, we invited artists who curated descriptive texts specific to landscape paintings. To increase the diversity of text descriptions, we provided multiple expressions for each text and specified the painting type within the descriptions. Finally, we saved the image paths and corresponding text descriptions in JSON files.
@@ 3.6 Evaluation Metrics
Learned Perceptual Image Patch Similarity. LPIPS is a metric to evaluate image similarity. It utilizes feature representations of deep learning models to measure the perceptual similarity between two images. Unlike traditional pixel-based similarity metrics, LPIPS focuses more on the perceptual quality of the image, that is, the perception of image similarity by the human visual system. The formula for LPIPS is given by:
@@
Gram Matrix. The gram matrix is a metric for evaluating the stylistic similarity of images and is commonly used in style migration tasks. It captures the texture information and stylistic features of an image by computing the inner product between the feature maps of a convolutional neural network. Specifically, the gram matrix describes the correlation between different channels in the feature map, thus reflecting the overall texture structure of the image.
@@
Histogram Similarity (Bhattacharyya Distance). The histogram similarity is a metric for evaluating the similarity of color distributions of two images and is widely used in image processing and computer vision. By comparing the color histograms of the images, the similarity of the images in terms of color can be determined.The bhattacharyya Distance  Bhattacharyya (1943) is a commonly used histogram similarity metric, especially for similarity calculation of probability distributions.
@@
Chamfer Match Score. The chamfer match score is a metric for evaluating the similarity of two sets of point clouds, widely used in computer vision and graphics. It quantifies the shape similarity between two sets of point clouds by calculating the average nearest distance between them. Specifically, chamfer match score is a variant of Chamfer Distance and is commonly used for tasks such as 3D shape matching, image alignment and contour alignment.
@@
Hausdorff Distance. The hausdorff distance is a metric for evaluating the maximum distance between two sets of point sets (point clouds, contours, etc.) and is widely used in computer vision, graphics, and pattern recognition. It measures the distance between the farthest corresponding points in two point sets, reflecting the degree to which they differ geometrically.
@@
Contour Match Score. The contour match score evaluates the similarity between the contour shapes of two images, making it particularly useful in applications where the overall shape and structure of objects are crucial. This score is determined by comparing the shape descriptors of the contours in the images.
@@
4 Experiment and Analysis
4.1 Generative Controllability
In order to prove that our proposed method can accurately control the structure and style of generated landscape images, we use different canny and style reference images as input to the LPGen model to test the effect of its generated images.
Figure 4: Diverse landscape paintings generated by our method. Our method, LPGen, is capable of producing artworks in various styles and creating differentiated paintings within the same style. Reference 1 through Reference 6 represent landscape paintings generated in the same style but with different canny edge maps. Canny 1 through Canny 6 depict landscape paintings generated with the same Canny edge map but using different style references.
As shown in Figure 4, we use the LPGen model to generate landscape paintings. The images in each column are landscapes generated by our proposed method in the same canny and different reference images. It is obvious that they can maintain the same structure and style as the reference image. For example, the reference image in Reference 3 is a azure green landscape style image. With different canny inputs, images with different structures are obtained. Obviously, its azure green landscape style is maintained. These experiments indicated that LPGen can learn the characteristics and styles of numerous classic landscape paintings and automatically create new highly realistic images. The generated landscape paintings feature typical natural elements such as mountains, rivers, trees, and stones and retain the ink color and brush techniques characteristic of traditional landscape paintings. By combining these two control methods, the LPGen model can generate highly realistic landscape paintings and flexibly adjust the style and details of the images, achieving precise control and innovative results in traditional art creation.
@@ 4.2 Qualitative Analysis
To accurately and objectively assess the effectiveness of the proposed LPGen model in generating high-quality landscape paintings, we conduct a comprehensive comparative analysis against several current state-of-the-art methods, including Reference Only Zhang et al. (2023), Double ControlNet  Zhang et al. (2023), and Lora  Hu et al. (2021).
Figure 5: Comparison between the proposed method and mainstream methods in generative landscape paintings. Figure 6a displays the constraint canny edge map, while Figure 6b shows the target ink-style reference image. The proposed method uses a Canny image to control the structure and a reference image to control the generated painting style. Each method generated landscape paintings using four different Canny edge maps, resulting in a total of 16 images.
Figure 5 (c) is generated by the Lora method. It has obvious noise and erroneous image information and cannot revert the style of the reference image. Compared with the Lora method, the image generated by the Reference Only method has a more accurate contour structure but introduces redundant contours, with indistinct style features and inaccurate colors, as shown in Figure 5 (d). The advantage of the Double ControlNet method is that the generated landscape paiting contain clear structure , but cannot effectively learn the style features of the reference image, as shown in Figure 5 (e).
Compared to other methods, the pros and cons of these methods are shown in Figure 5, which shows that the method proposed in this study was the best among all tested state-of-the-art methods. The proposed LPGen effectively addresses significant issues such as poor style transfer and blurred lines in generated images, resulting in high-quality landscape paintings. LPGen not only preserves the style and structure of the target photos but also successfully captures the essence of the ink wash style, thereby achieving superior overall quality and fidelity.
4.3 Quantitative analysis
Table 2: Quantitative comparison of the proposed LPGen with several state-of-the-art models. LPIPS evaluates perceptual similarity, gram matrix (GM) measures texture correlations for style similarity, and histogram similarity (HS) assesses color distribution. chamfer match score (CMS i) focuses on average edge similarity, hausdorff distance (HD) emphasizes worst-case edge similarity, and contour match score (CMS ii) evaluates overall shape similarity.
@@ 4.3 Quantitative analysis
The landscape paintings generated by the proposed method were quantitatively compared with those generated by Reference Only, Double ControlNet, and Lora. Each model generated 15 images for each pair of style and structure, resulting in a total of 60 landscape painting images. For each model, the highest values for six metrics of the generated images were recorded: gram matrix similarity, histogram similarity, LPIPS, chamfer match score, hausdorff distance, and contour match score. The scores of the different diffusion models are shown in Table 2.
@@
In Table 2, the first three metrics, the gram matrix, histogram similarity, and LPIPS, are used to analyze the structural similarity between model outputs and reference images. The table shows that LPGen performs best with respect to the gram matrix, suggesting that LPGen’s generated images have the most similar texture style to the reference image. Furthermore, LPGen performs best with respect to histogram similarity, suggesting that LPGen-generated images have the most similar color distribution to the reference image. For structural similarity, LPGen excels in both gram matrix and histogram similarity, while Lora performs best in LPIPS due to the structure controller. To repetitive fine-tuning, Lora performs best in terms of LPIPS, indicating that Lora’s generated images are most perceptually similar to the reference image. Taking these factors into account, LPGen is the best model for overall structural Tsimilarity.
@@
In Table 2, the last three metrics, the chamfer match score, hausdorff distance, and contour match score, are used to analyze the style similarity between different model outputs and the reference image, mainly focusing on edges and contours. Obviously, LPGen performs best with respect to the chamfer match score, suggesting that LPGen’s generated images have the most similar edges to the reference image. Meanwhile, LPGen performs best with respect to the hausdorff distance, suggesting that LPGen’s generated images are closest to the reference image regarding edge similarity. Furthermore, LPGen performs best with respect to the contour match score, indicating that LPGen’s generated images have the most similar contour shapes to the reference image. Considered comprehensively, the landscape paintings generated by the proposed model have the most similar edges and contour shapes to the reference image, making it the best model for style similarity.
4.4 Visual Assessment